OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $1,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain ho

OpenStudy (anonymous):

Please explain how you arrived at your answer and show all of your work.

OpenStudy (anonymous):

Ok let's do option 1 first...if you place 1000$ down, how much do you have left to pay?

OpenStudy (anonymous):

15,000?

OpenStudy (anonymous):

Correct. Now spread that out over 60 months...how much are you paying per month?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

Ok good so thats the monthly amount for option 1

OpenStudy (anonymous):

For option 2, we should first figure out the cost of the car as original cost*(1.02)*6

OpenStudy (anonymous):

16000*0.02*72 = 23040 23040/72 = 320

OpenStudy (anonymous):

isint it supposed to be 72 months

OpenStudy (anonymous):

not 6

OpenStudy (anonymous):

but it's an annual interest rate right 72 months = 6 years

OpenStudy (anonymous):

oooohhh

OpenStudy (anonymous):

ok wat about the 1.02?

OpenStudy (anonymous):

where did that come from?

OpenStudy (anonymous):

So .02 is the interest rate but you still have to add that to the original price. So it's original price+.02*original price*6. I just factored out original price.

OpenStudy (anonymous):

If it makes more sense just calculate the final cost as FC = OP+.02*OP*6

OpenStudy (anonymous):

16,000 * 0.02 * 6 = 1920

OpenStudy (anonymous):

1920÷72 = $26?

OpenStudy (anonymous):

1920 is the interest you have to pay on the car...but remember you have to pay the original price of 16,000 also. the total price is 17920 right

OpenStudy (anonymous):

ok so is this wat i have to do 16,000 • 0.02 • 6 = 1920 1920 + 16000 = 17920 17920 ÷ 72 = $248.8

OpenStudy (anonymous):

Correct. Now just pick which option is lower and bingo

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (anonymous):

Np